OpenStudy (anonymous):

I need help really bad on pythagorean theorem!!!!!For a right triangle that has legs a and b and hypotenuse c, find the length of the side not given here. a=1 c=√3

OpenStudy (chillout):

Remember that a²+b² = c² and plug in the values.

OpenStudy (jhannybean):

\[\large a^2+b^2=c^2\]\[\large b^2= c^2-a^2\] use givens.\[\large b^2=(\sqrt{3})^2-(1)^2\]\[\large b^2=3-1\]\[\large b=\sqrt{3-1}\]i'll letyou finish up.